OpenStudy (anonymous):

Which of the following is the solution to -x - 10 = -3? A. x = -13 B. x = 13 C. x = 7 D. x = -7

OpenStudy (anonymous):

D

OpenStudy (raden_zaikaria):

point me ur step... i will check for u... k

OpenStudy (anonymous):

letter c as x-10=-3 x=10-3 x=7

OpenStudy (anonymous):

add 10 to each side making -x=7

OpenStudy (anonymous):

then switch the signs on both sides, making x=-7

OpenStudy (anonymous):

ya i got 7 too ! i just plugged the number in for x and that was my answer ! thanks ! (:

OpenStudy (anonymous):

therefore, the answer is D, x=-7

OpenStudy (anonymous):

that was wrong i didnt saw the negative sign of x the correct one is c -x-10=-3 -x=10-3,transpose 10 beside 3 and the negative sign of 10 will become positive as it passes through the - sign. -x=7 x=7divide no.7 by the negative sign of x

OpenStudy (raden_zaikaria):

answer is -7 because..... -x not x

OpenStudy (raden_zaikaria):